OpenStudy (anonymous):

Take the second derivative of y= sqrt(x^2+1)

OpenStudy (nincompoop):

take the first derivative

OpenStudy (anonymous):

Yes, and I got y'= x(x^2+1)^-1/2

OpenStudy (nincompoop):

then take the derivative of that

OpenStudy (anonymous):

Do I use the product rule of second derivative?

OpenStudy (nincompoop):

how would you normally solve that?

OpenStudy (anonymous):

I got y " = -x^2(x^2+1)^-3/2

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

@nincompoop ??

OpenStudy (rational):

Alternative : \[y = \sqrt{x^2+1} \implies y^2 = x^2+1\] implicitly differentiating both sides gives \[2yy' = 2x \implies yy' = x\] differentiate again implicitly and solve \(y'\)

OpenStudy (michele_laino):

hint: you have to apply the derivative of a quotient between two functions, namely: \[\Large y'' = \frac{{1 \cdot \sqrt {{x^2} + 1} - x \cdot \frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}} = ...?\]

OpenStudy (michele_laino):

since, the first derivative, is: \[\Large y' = \frac{x}{{\sqrt {{x^2} + 1} }}\]

OpenStudy (nincompoop):

ye you can use quotient rule for that, but you can also use product rule with with chain rule

OpenStudy (nincompoop):

it depends on how you can strategize and make it simpler for you to solve

OpenStudy (anonymous):

I tried usinf the product rule and chain rule, but I kept on getting y '' = -x^2(x^2+1)^-3/2

OpenStudy (nincompoop):

write the product rule for me

OpenStudy (rational):

Easy to make mistakes with those messy expressions. Just take a deep breath and double check ur work

OpenStudy (nincompoop):

laughing inside my head and physically smiling

OpenStudy (michele_laino):

are you sure? @esam2

OpenStudy (anonymous):

I got this for using the product rule and chain rule: \[y''= (x^2+1)^{-1/2}+x(-1/2(x^2+1)^{-3/2}(2x)\]

OpenStudy (anonymous):

@rational @Michele_Laino

OpenStudy (michele_laino):

using the quotient rule, I got this: \[\Large \begin{gathered} y'' = \frac{{1 \cdot \sqrt {{x^2} + 1} - x \cdot \frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}} = \frac{{{x^2} + 1 - {x^2}}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} = \hfill \\ \hfill \\ = \frac{1}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

That is the right answer, I'm just wondering how to do it by using the product rule...

OpenStudy (michele_laino):

then, we have to write the first derivative, as below: \[y' = \frac{x}{{\sqrt {{x^2} + 1} }} = x \cdot \frac{1}{{\sqrt {{x^2} + 1} }} = x{\left( {{x^2} + 1} \right)^{ - 1/2}}\]

OpenStudy (anonymous):

Yes, I got that part

OpenStudy (michele_laino):

so, we have: \[y'' = 1 \cdot {\left( {{x^2} + 1} \right)^{ - 1/2}} + x\left( { - \frac{1}{2}{{\left( {{x^2} + 1} \right)}^{ - 3/2}}2x} \right) = ...?\]

OpenStudy (michele_laino):

\[\begin{gathered} y'' = 1 \cdot {\left( {{x^2} + 1} \right)^{ - 1/2}} + x\left( { - \frac{1}{2}{{\left( {{x^2} + 1} \right)}^{ - 3/2}}2x} \right) = \hfill \\ = \frac{1}{{\sqrt {{x^2} + 1} }} - \frac{{{x^2}}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} = \frac{{{x^2} + 1 - {x^2}}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

How did you get it like that?

OpenStudy (michele_laino):

I got that, using the properties of the real numbers

OpenStudy (anonymous):

What is the properties of the real numbers?

OpenStudy (michele_laino):

for example, what is the least common multiple between: \[{\sqrt {{x^2} + 1} }\] and \[{{{\left( {{x^2} + 1} \right)}^{3/2}}}\]?

OpenStudy (anonymous):

I didn't quite understand how to get this part:\[\frac{ x^2+1+x^2 }{ (x^2+1)^{3/2} }\]

OpenStudy (anonymous):

sqrt(x^2+1)?

OpenStudy (michele_laino):

no, it is: \[{{{\left( {{x^2} + 1} \right)}^{3/2}}}\]

OpenStudy (anonymous):

Oh..

OpenStudy (michele_laino):

so, we can write: \[\large \frac{1}{{\sqrt {{x^2} + 1} }} - \frac{{{x^2}}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} = \frac{{1 \cdot \left( {{x^2} + 1} \right) - {x^2} \cdot 1}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}}\]

OpenStudy (anonymous):

Oh I see!

OpenStudy (anonymous):

I think I got it from here. Thanks for all your help! :)

OpenStudy (michele_laino):

Thanks! :)